1. Calculate the efficiency of a steam engine working at a temperature 573 K assuming the temperature of the surrounding as 298 K.

2. How much useful work can be done by carnot engine working between 0°C and 100°C if the heat supplied is 1900kJ.

3. Calculate the amount of heat supplied to a carnot cycle working between 368K and 288K if maximum obtain is 895J.
1
Expert's answer
2020-05-01T14:30:29-0400

 emax = (Tsteam- Tair)/Tsteam = "\frac{573-298}{573}=0.48" or 48%


efficiency of carnot engine = \"1-\frac{T_L}{T_H}=1-\frac{273}{373}=0.27"

useful work = heat "\times" efficiency = 1900"\times"0.27=513 J